OpenStudy (anonymous):

Which expression is equivalent to 11(17 – x)? A. 11 • 17 – 11x B. 11 • 17 – x C. 11 + 17 – x D. 17 – 11x

OpenStudy (kidrah69):

What do you think? (hint think of how you would solve that)

OpenStudy (kidrah69):

Think of the distributive property steps...

OpenStudy (anonymous):

I was thinking B

OpenStudy (kidrah69):

Thats not correct

OpenStudy (skullpatrol):

Why?

OpenStudy (kidrah69):

?

OpenStudy (kidrah69):

When you distribute or i should say solve the equation....you will not end up with that. We are missing something....do you know what that is?

OpenStudy (skullpatrol):

$$\Huge a(b+c)=ab+ac$$

OpenStudy (anonymous):

No...

OpenStudy (kidrah69):

Do you know how to solve that? the main equation? ( i know we are not supposed to solve it but it will help you understand)

OpenStudy (skullpatrol):

Notice how the "a" is distributed over the b and c

OpenStudy (anonymous):

Yes, I solve the parentheses first then multiply right?

OpenStudy (kidrah69):

No...always distribute the outter number into the parenthesis.....this is known as the distributive property.

OpenStudy (skullpatrol):

This is called an EXPRESSION not an equation ... You can only SOLVE equations :-)

OpenStudy (anonymous):

Oh okay lol I am so confused :3

OpenStudy (kidrah69):

Right what i was trying to get at was if she knows how to distribute then she can get her answer....

OpenStudy (kidrah69):

@~*Ashtyn*~ i think i'lll let @skullpatrol finish i dont want to confuse you :)

OpenStudy (skullpatrol):

Look carefully at the distributive property of multiplication with respect to subtraction.

OpenStudy (skullpatrol):

$$\Huge a(b-c)=ab-ac$$

OpenStudy (anonymous):

Okay...

OpenStudy (skullpatrol):

Replace a with 11 in your question

OpenStudy (skullpatrol):

$$\Huge 11(17-x)=?$$

OpenStudy (skullpatrol):

b is 17 and c is x

OpenStudy (anonymous):

Sorry my acc. lagged and logged itself out so just give me one minute :)

OpenStudy (skullpatrol):

$$\Huge a(b-c)=ab-ac$$ $$\Huge 11(17-x)=?$$

OpenStudy (anonymous):

\[11\times17-17\times x\]

OpenStudy (anonymous):

So A

OpenStudy (skullpatrol):

Correct :D

rvc (rvc):

candy(2child-3)=2candy(child)-3candy

rvc (rvc):

All the best

OpenStudy (anonymous):

Combine like terms. 8 + 6x – 11x + 12 A. 14x + 1x B. 20 – 5x C. 20 – 17x <<<<<< D. 20 + 5x

OpenStudy (skullpatrol):

What is 6-11?

OpenStudy (anonymous):

5

OpenStudy (skullpatrol):

$$\Huge 6-11=6+(-11)$$

OpenStudy (anonymous):

???

OpenStudy (anonymous):

It says combine like terms... Wouldn't B be correct?...

OpenStudy (skullpatrol):

Yes

OpenStudy (anonymous):

I mean C wouldn't C be correct lol

OpenStudy (skullpatrol):

B is the right answer.

OpenStudy (skullpatrol):

$$\Huge 6-11 =6+(-11)=-5$$

OpenStudy (skullpatrol):

$$\Huge 6x-11x=(6-11)x=-5x$$

OpenStudy (skullpatrol):

$$\Huge ba-ca=(b-c)a$$

OpenStudy (skullpatrol):

That^ is the distributive property

OpenStudy (anonymous):

I'm combining like terms :3

OpenStudy (anonymous):

I thought......

OpenStudy (skullpatrol):

What do you think the definition of "like terms" is ?

OpenStudy (skullpatrol):

"Like" terms are also called "similar" terms

OpenStudy (skullpatrol):

They are the same except for their numerical coefficient

OpenStudy (skullpatrol):

6x and 11x are like terms because they are the same except for the numerical coefficients of 6 and 11, right?

OpenStudy (skullpatrol):

They have a common factor of x.

OpenStudy (anonymous):

I get it lol I picked B